OpenStudy (anonymous):

At a science competition, 25 problems were given. Each correct answer earned 4 points, and 2 points were deducted for each incorrect answer. No points were deducted for any unanswered problem. Logan answered all but one of the problems, receiving a score of 66. How many correct answers did Logan have?

OpenStudy (anonymous):

please help i will give a medal

OpenStudy (anonymous):

please help

ganeshie8 (ganeshie8):

So Logan answered 24 problems

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now we have to tell how many correct he answered

ganeshie8 (ganeshie8):

Out of 24 problems that he answered, `x` problems were correct and `y` problems wrong

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

do u know how many

OpenStudy (anonymous):

at last he got 66 points

ganeshie8 (ganeshie8):

can we setup below equations : x + y = 24 4x-2y = 66 ?

OpenStudy (anonymous):

how to do that

OpenStudy (anonymous):

he answered 24 problems so the maximum score he should get is 94 because 24 x 4 is 96

OpenStudy (anonymous):

let me give u a medal first u gave time and i am really thankfull to u

ganeshie8 (ganeshie8):

you're right, solve above equations, "x" is the number of correct answers

OpenStudy (anonymous):

can u solve that for me i am weak in math thank u

ganeshie8 (ganeshie8):

x + y = 24 4x-2y = 66 from first equation, y = 24-x plug that in second equaiton : 4x - 2(24-x) = 66 4x - 48 + 2x = 66 6x = 114 x = 19

ganeshie8 (ganeshie8):

trying them on your own makes you strong :)

OpenStudy (anonymous):

thank u a lot

OpenStudy (anonymous):

i dont know what to say you are the best person i have seen in world

ganeshie8 (ganeshie8):

you're welcome ! so he answered 19 questions correctly, 5 questions incorrectly, and 1 question left

OpenStudy (anonymous):

okay thank u a lot

ganeshie8 (ganeshie8):

19*4 = 76 5*2 = 10 so his overall score is 76-10 = 66 as it should be, so we're good

OpenStudy (anonymous):

do u remember a member named The.Great.Stupid.PErson

OpenStudy (anonymous):

@ganeshie8 do u remember a member whose name was The.Great.Stupid.Person